# Empirical validation of the polarization transition in a double-random field model of elections

### v3

@@ Abstract
Calibration to US House election data for Republican-leaning races ($p=0.6\pm 0.05)$. To illustrate the effectiveness of the classification model also on the case of a subset of races, we choose the races whose previous results were in the range corresponding to $p=0.6\pm 0.05$. The subset consists of 1145 election races. We find the optimal classification model by finding $T$ and $h_{c}$ that maximize accuracy. We find out that while the optimal parameters slightly change ($T^{\star}=0.845$, $h_{c}=2$ million USD), the overall behavior does not change. We should also stress that since the smaller size of the subset, the statistical tests like the McNemar test exhibit a bit weaker (but still significant) value, which is caused by the fact that the number of points that are classified differently by the optimal model ($T=T^{\star}$) and the null model ($T=1$) is relatively low.
Comparison of calibration of US House election data for different time periods.
@@
Figure 8: Comparison of fitted parameters for US House elections in four decades corresponding to the previous plots, focused on incumbency regions
@@
Comparison of calibration of US House election data with a support vector machine model. Finally, we compare our method with a standard machine learning classification method, particularly the support vector machine. Again, we focus on the close races, i.e., all races with $p=0.5\pm 0.05$. In order to utilize the natural symmetry of the system, i.e., $h^{+}=h^{-}$ leading to $m=0$, we transform the spending data into the following features:
$\sigma_{h}$ $=$ $\frac{{h^{+}}+{h^{-}}}{2}$
$\delta_{h}$ $=$ $h^{+}-h^{-}$
@@
$h^{+}$ $=$ $\sigma_{h}+\delta_{h}/2$
$h^{-}$ $=$ $\sigma_{h}-\delta_{h}/2$
We transform the classifier into the original space. By measuring the intersection of the incumbency region with the diagonal ($h^{+}=h^{-}$), we obtain the estimate on the equivalent of the critical threshold, which is here $h^{*}=\mathdollar 2.68M$. This threshold is higher than predicted by the model in the main text. Furthermore, the accuracy on the whole dataset is lower (only $80\%$), possibly due to overfitting.
Figure 9: Application of SVM to estimate the incumbency region.

### v4

@@ Abstract
Calibration to US House election data for Republican-leaning races ($p=0.6\pm 0.05)$
To illustrate the effectiveness of the classification model also on the case of a subset of races, we choose the races whose previous results were in the range corresponding to $p=0.6\pm 0.05$. The subset consists of 1145 election races. We find the optimal classification model by finding $T$ and $h_{c}$ that maximize accuracy. We find out that while the optimal parameters slightly change ($T^{\star}=0.845$, $h_{c}=2$ million USD), the overall behavior does not change. We should also stress that since the smaller size of the subset, the statistical tests like the McNemar test exhibit a bit weaker (but still significant) value, which is caused by the fact that the number of points that are classified differently by the optimal model ($T=T^{\star}$) and the null model ($T=1$) is relatively low.
Figure 8: Comparison of fitted parameters for US House elections in four decades corresponding to the previous plots, focused on incumbency regions.
Comparison of calibration of US House election data for different time periods
@@
Figure 9: Application of SVM to estimate the incumbency region.
Comparison of calibration of US House election data with a support vector machine model
Finally, we compare our method with a standard machine learning classification method, particularly the support vector machine. Again, we focus on the close races, i.e., all races with $p=0.5\pm 0.05$. In order to utilize the natural symmetry of the system, i.e., $h^{+}=h^{-}$ leading to $m=0$, we transform the spending data into the following features:
$\sigma_{h}$ $=$ $\frac{{h^{+}}+{h^{-}}}{2}\,$,
$\delta_{h}$ $=$ $h^{+}-h^{-}\,.$
@@
$h^{+}$ $=$ $\sigma_{h}+\delta_{h}/2\,$,
$h^{-}$ $=$ $\sigma_{h}-\delta_{h}/2\,.$
We transform the classifier into the original space. The fitted region is depicted in Fig. 9. By measuring the intersection of the incumbency region with the diagonal ($h^{+}=h^{-}$), we obtain the estimate for the equivalent of the critical threshold, which is here $h^{*}=\mathdollar 2.68M$. This threshold is higher than predicted by the model in the main text. Furthermore, the accuracy on the whole dataset is lower (only $80\%$), possibly due to overfitting.